### v2

# Structured Probabilistic Inference and Generative Modeling

Abstract
Understanding causality in event sequences where outcome labels such as diseases or system failures arise from preceding events like symptoms or error codes is critical—yet remains an unsolved challenge across domains like healthcare or vehicle diagnostics. We introduce CARGO, a scalable multi-label causal discovery method for sparse, high-dimensional event sequences comprising of thousands of unique event types. Using two pretrained causal Transformers as domain-specific foundation models for event sequences—CARGO infers in parallel, per sequence one-shot causal graphs and aggregates them using an adaptive frequency fusion to reconstruct the global Markov boundaries of labels. This two-stage approach enables efficient probabilistic reasoning at scale while bypassing the intractable cost of full-dataset conditional independence testing. Our results on a challenging real-world automotive fault prediction dataset with over 29,100 unique event types and 474 imbalanced labels demonstrate CARGO’s ability to perform structured reasoning.
1 Introduction
Understanding why specific events lead to particular outcomes is vital for effective diagnosis, predictions and overall decision making [liu2025learning, shtp]. For instance, ”what serie of events captured by diagnostic led to this vehicle failure” or ”what symptoms led to this disease” [MANOCCHIO2024122564, MedBERT, bihealth, flight_service_cd, pdm_dtc_feature_extraction]. Here, an event sequence consists of a list of discrete events $x_{i}$ recorded asynchronously over time, while labels $\boldsymbol{y}$ summarize outcomes associated with the full sequence (e.g, a diagnosed defect or condition).
A fundamental obstacle in these settings is dimensionality. Real-world systems often involve tens of thousands of possible events, which renders the causal discovery intractable for current algorithms [hasan2023a]. To address this, we reinterpret multi-label causal discovery for event sequences as a form of Bayesian model averaging [HOET1999, pearl_1998_bn], where each sequence is treated as a sample from a local causal model. Specifically, each sequence induces a one-shot causal graph (i.e., a directed acyclic graph (DAG)) [tf_causalinterpretation_neurips_2023]. Together they can be fused to form a unified global structure [divide_conquer_mb_discovery]. This process, known as structural fusion [consensus_bn_jose], aggregates local graphs into a consensus causal graph over all observed sequences.
To summarise our contributions: we introduce CARGO (Causal Aggregation via Regressive Graph Operations), the first method to provide scalable causal discovery across thousands of labelled event sequences, with theoretical guarantees under standard assumptions. It is divided into two phases: (1) One-shot graph extraction, where for each sequence CARGO infers the local Markov boundary of each label using two autoregressive Transformers as density estimators [tf, density_estimator_2O21_journal_ci] (2) Graph fusion, where the local graphs are aggregated via an adaptive thresholding function to provide global Markov Boundaries.
We empirically validate CARGO on a large-scale vehicular dataset composed of about $29,100$ events and $474$ imbalanced labels, demonstrating for the first time, scalability and practical superiority over traditional causal discovery baselines. We also provide ablation on scoring criterions, frequency thresholds and the quality of the Transformers.
2 Preliminary & Related Work
A full description of the notations and definitions used throughout the paper can be found in Appendix A.
Event Sequence Modelling. Event sequences are typically represented as a series of time-stamped discrete events $S=\{(t_{1},x_{1}),\ldots,(t_{L},x_{L})\}$ where $0\leq t_{1}<\ldots\leq t_{L}$ denotes the time of occurrence of event type $x_{i}\in\mathbb{X}$ drawn from a finite vocabulary $\mathbb{X}$. In multi-label settings, a binary label vector $\boldsymbol{y}\in\{0,1\}^{|\mathbb{Y}|}$ is attached to $S$ and indicates the presence of multiple outcome labels chosen from $\mathbb{Y}$ occuring at the final time step $t_{L}$. Together, this results in a multi-labeled sequence $S_{l}=(S,(\boldsymbol{y}_{L},t_{L}))$.
Event sequence modelling has been widely applied to predictive tasks. For instance, in the automotive domain, Diagnostic Trouble Codes (DTCs) [pdm_dtc_feature_extraction] are logged asynchronously over time and used to infer failures or error patterns [math2024harnessingeventsensorydata]. In healthcare, electronic health records encode temporal sequences of symptoms to perform predictive tasks [MedBERT, pmlr-v219-labach23a, bihealth]. A common modelling strategy [crf, cmm] separates such event types $\mathbb{X}$ from labels $\mathbb{Y}$.
Transformers [tf, gpt, touvron2023llamaopenefficientfoundation] have emerged as the dominant architecture for sequence modelling. Recent work leveraged Transformers in high-dimensional event spaces for next-event and label prediction. math2024harnessingeventsensorydata proposed a dual Transformer architecture where one model predicts the next event type (DTC), and the other predicts label occurrence (e.g, error pattern). Through this paper, we repurpose this dual architecture for causal discovery.
Multi-label Causal Discovery seeks to identify the Markov Boundary (MB) of each label—its minimal set of parents, children, and spouses—such that the label is conditionally independent of all other variables given its MB [optimal_feature_set_cd] (Def. 3).
While classical constraint-based algorithms have shown success on low-dimensional tabular data [constrainct_based_cd, causality_based_feature_selection_2019], their application to event sequences with multi-label outputs remains challenging due to: (1) dimensionality—thousands of event types increase super-exponentially the number of graphs (2) sparsity—multi-hot encodings often underrepresent rare but important events (3) distributional assumptions—such as linearity or Gaussian noise, which rarely hold in real-world sequences [cd_temporaldata_review].
Contemporary works point toward dividing classical causal discovery for high-dimensional datasets into sub-problems and graph aggregation. laborda_ring_based_distirbuted introduce a ring-based distributed algorithm for learning high-dimensional BN, divide_conquer_mb_discovery explores a distributed approach for large-scale causal structure learning and recursive_mb_approach_pmlr_large_scale for Markov Boundaries.
Bayesian Network [pearl_1998_bn] has served as a modelling technique for a variety of decision problems. It is defined as a triplet $\boldsymbol{U},\mathbb{G},P>$ with $P$ the joint distribution over a variable set $\boldsymbol{U}$ of a directed acyclic graph $\mathbb{G}=(\boldsymbol{U},E)$ with $E$ as the set of directed edges. This triplet must satisfy the Markov Condition: every random variable $U_{i}$ is independent of its non-descendant variables given its parents $\text{Pa}(U_{i})$ in $\mathbb{G}$. The directed edge $(U_{i}\rightarrow U_{j})$ encodes a probabilistic dependence. Thus, the joint probability distribution can be factorised as:
$P(U_{1},\cdots,U_{n})=\prod^{n}_{i=1}P(U_{i}|\text{Pa}(U_{i}))$
The DAG encodes a set of conditional independencies $\mathcal{I}(\mathbb{G})$, where each element corresponds to a conditional independence relation $U_{i}\perp U_{j}|\boldsymbol{Z}$, meaning that $U_{i}$ and $U_{j}$ are conditionally independent given the set of variables $\boldsymbol{Z}$. Formally, a DAG $\mathbb{G}_{k}$ is an I-map (or Independence map) of another $\mathbb{G}$ if the set of conditional independencies encoded by $\mathbb{G}_{k}$ is a subset of those encoded by $\mathbb{G}$:
$\mathcal{I}(\mathbb{G}_{k})\subseteq\mathcal{I}(\mathbb{G})$
[…]
Figure 1: An example of a causal graph extracted from a single multi-label event sequence where $\text{MB}_{1}$ represents the Markov Boundary of $Y_{1}$ and $\text{MB}_{2}$ the Markov Boundary of $Y_{2}$.
[…]
Bayesian Model Averaging. Fusing BN has several direct applications. Either to average multiple models from different experts to learn a global and average representation [HOET1999]. Or to perform causal discovery in distributed settings with federated learning algorithms [fedpc_fedfci, Guo_Yu_Liu_Li_2024].
Formally, Given a set of Bayesian Networks $\{B_{k}\}^{m}_{k=1}$ with associated DAGs $\{\mathbb{G}_{k}=(V_{k},E_{k})\}^{m}_{k=1},V_{k}\in\boldsymbol{U}$ sharing the same finite set of node $\boldsymbol{U}$, structural fusion aim to construct the DAG $\mathbb{G}^{*}=(V,E),V\in\boldsymbol{U}$. Multiple fusion methods exist and leverage either the probability distribution $p$ by doing Bayesian Model Averaging [HOET1999] or focus on the structural learning (Fig. 8) of $\mathbb{G}^{*}$ [delSagrado2001, consensus_bn_jose, go2022robust, approx_fusionPUERTA2021155, Guo_Yu_Liu_Li_2024], which is an NP-hard [consensus_bn_jose] problem.
We will focus on the second element in this paper. Hence, we are seeking the merged edges $E=\bigcup^{m}_{i=1}E^{\sigma}_{i}$. The consistent node ordering $\sigma$ ensures acyclicity. The fused DAG $\mathbb{G}^{*}$ is the minimal I-map of the intersection of the conditional independencies across all DAGs $\mathbb{G}_{k}=(V_{k},E_{k})\}^{m}_{k=1}$.
Greedy Equivalence Search. (GES) [ges] is one of the most theoretically sound methods to recover a Markov equivalence class (MEC, Def. 4) of a DAG. Particularly in the context of large samples of data, GES provides theoretical guarantees of reaching the true graph. Formally, GES searches for the MEC of the graph $\mathbb{G}^{*}$ from the observational dataset $\boldsymbol{D}$ with distribution $p$. This defines the optimisation problem as:
$\mathbb{G}^{*}=\arg_{\mathbb{G}}\text{max}\;S(\mathbb{G},\boldsymbol{D})$ (1)
ges proved that under parametric assumption, a large number of samples and using the Bayesian Information Criterion (BIC) as criterion $S$, GES is guaranteed to recover a Markov Equivalence Class of $\mathbb{G}^{*}$.
Frequency Based. Multiple heuristics have been developed to merge multiple BNs. One is edge frequency cutoff [Steele2008Consensus], other on estimating the proportion of false positive edges [consensus_bn_freq_cutoff_no] and integer linear programming (ILP) [consensus_bn_freq_cutoff_no] or a mix of both: [torrijos2025informedgreedyalgorithmscalable]. As pointed out by [consensus_bn_freq_cutoff_no], choosing a frequency cutoff $\tau$ is particularly challenging. Moreover, under class imbalance and long-tail problem in classification [longtailissue], the theoretical property of frequency approaches, such as the law of large numbers, might not hold.
We will now explain the two phases of CARGO, which are (1) One-shot causal discovery (2) Graph aggregation using adaptive thresholding. The Proofs of Lemmas and Theorems can be found in the Appendix B.
3 One-Shot Causal Discovery
Let $S^{k}_{l}$ be a multi-labeled sequence drawn from a dataset $D=\{S^{1}_{l},\cdots,S^{m}_{l}\}\subset\mathbb{S}$ and $\mathbb{G}_{k}$ the sequential BN (Fig. 1) with attached labels. The goal of multi-label causal discovery is to identify the Markov Boundary of each label $Y_{j}\in\boldsymbol{Y}$ present in $S^{k}_{l}$. To be able to access conditional independence (Def. 2) between $X_{i}$ and $Y_{j}$ conditioned on the past events $\boldsymbol{Z}=(x_{1},\cdots,x_{i-1})=S_{<i}$, we model the event apparitions using a sequential BN (Fig. 1).
We would like to assess how much additional information event $X_{i}$ occurring at step $i$ provides about label $Y_{j}$ when we already know the past sequence of events $\boldsymbol{Z}=S_{<i}$. We essentially try to answer if:
$P(Y_{j}|X_{i},\boldsymbol{Z})=P(Y_{j}|\boldsymbol{Z})\Leftrightarrow D_{KL}(P(Y_{j}|X_{i},\boldsymbol{Z})\|P(Y_{j}|\boldsymbol{Z}))=0$
where $D_{KL}$ denotes the Kullback-Leibler divergence [cover1999elements]. The distributional difference between the conditionals $P(Y_{j}|X_{i},\boldsymbol{Z}),P(Y_{j}|\boldsymbol{Z})$ is akin to Information Gain $I_{G}$ [quinlan:induction] conditioned on past events:
$I_{G}(Y_{j},x_{i}|z)\triangleq D_{KL}(P(Y_{j}|X_{i}=x_{i},\boldsymbol{Z}=z))||P(Y_{j}|\boldsymbol{Z}=z))$ (2)
Which is equals to the difference between the conditional entropies [cover1999elements] denoted as $H$:
$I_{G}(Y_{j},x_{i}|z)=H(Y_{j}|z)-H(Y_{j}|x_{i},z)$ (3)
More generally, we can access the conditional independence of event $X_{i}$ and label $Y_{j}$ using the conditional mutual information (CMI) [cover1999elements] which is simply the expected value over $z$ of the information gain $I_{G}(Y_{j},X_{i}|z)$ such as:
$I(Y_{j},X_{i}|\boldsymbol{Z})\triangleq H(Y_{j}|\boldsymbol{Z})-H(Y_{j}|\boldsymbol{Z},X_{i})=\mathbb{E}_{p(z)}[I_{G}(Y_{j},X_{i}|\boldsymbol{Z}=z)]$ (4)
It can be interpreted as the expected value over all possible contexts $\boldsymbol{Z}$ of the deviation from independence of $X_{i},Y_{j}$ in this context. To approximate Eq. (4), a naive Monte Carlo [Doucet2001] approximation is performed where we draw $N$ random variations of the conditioning set $z^{(l)}=\{x^{(l)}_{0},\ldots,x^{(l)}_{i-1}\}$, denoting the $l$-th sampled particle:
$\hat{I}_{N}(Y_{j},X_{i}\mid\boldsymbol{Z})$ $=\frac{1}{N}\sum_{l=1}^{N}I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z^{(l)})$ (5)
This estimator is unbiased because the contexts $z^{(l)}$ are sampled directly from $\text{Tf}_{x}$ using a proposal $Q$ with the same support as $P(\boldsymbol{Z})$. Since $I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z)$ is a difference between conditional entropies ((2)), it is thus bounded uniformly [cover1999elements] by the log of supports such as:
$0<I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z^{(l)})=H(Y_{j}|z^{(l)})-H(Y_{j}|X_{i},z^{(l)}))\leq H(Y_{j})\leq\log{|\mathbb{Y}|}$
Thus the posterior variance of $f_{i}=I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z^{(l)})$ satisfies $\sigma^{2}_{f_{i}}\triangleq\mathbb{E}_{p(z)}[f^{2}_{i}(p(z)]-I^{2}(f_{i})<+\infty$ [Doucet2001] then the variance of $\hat{I}_{N}(f_{i})$) is equal to $\textit{var}(\hat{I}_{N}(f_{i}))=\frac{\sigma^{2}_{f_{t}}}{N}$ and from the strong law of large numbers:
$\hat{I}_{N}$ $\xrightarrow[N\to+\infty]{\text{a.s.}}\mathbb{E}_{p(z)}\!\left[I_{G}(Y_{j},X_{i}\mid\boldsymbol{Z}=z)\right]\triangleq I(f_{i}).$ (6)
Density Estimation. We used two pretrained Transformers $(\text{Tf}_{x},\text{Tf}_{y})$ trained via maximum likelihood on a dataset of multi-labelled event sequences $D=\{S^{1}_{l},\cdots,S^{m}_{l}\}\subset\mathbb{S}$. We assume that they perfectly model the true conditional distributions of events and labels (A4). While due to the strict equivalences between CI-tests and conditional independence, it is difficult to provide a theoretical guarantee under imperfect models, we acknowledge that this assumption may be violated. We note that most of the causal discovery methods either have strong parametric data assumptions or assume a perfect CI-test. We provide an ablation study on the impact of the NADE’s quality on the one-shot phase in Appendix C.1, including number of parameters, context $c$ and $\text{Tf}_{y}$ performance.
Formally, the two Transformers infer the probability of the next event and label conditioned on the past events using the hidden states $\boldsymbol{h}^{x}_{i-1},\boldsymbol{h}^{y}_{i}\in\mathbb{R}^{d}$, from $\text{Tf}_{x},\text{Tf}_{y}$ respectively:
$\text{Tf}_{x}(S_{<i})$ $=\textit{Softmax}(\boldsymbol{h}^{x}_{i-1})=P_{\theta_{x}}(X_{i}|\boldsymbol{Z}),\;\hat{x}^{(z)}_{i}\triangleq P_{\theta_{x}}(X_{i}=x_{i}|\boldsymbol{Z})$ (7)
$\text{Tf}_{y}(S_{\leq i})$ $=\textit{Sigmoid}(\boldsymbol{h}^{y}_{i})=P_{\theta_{y}}(Y|X_{i},\boldsymbol{Z}),\hat{y}^{(z)}_{i+1,j}\triangleq P_{\theta_{y}}(Y=y_{j}|X_{i}=x_{i},\boldsymbol{Z})$ (8)
Sequential One-shot Causal Discovery. The CMI using Eq.(4) is computable only with the posteriors $P(Y_{j}|\boldsymbol{Z}),P(Y_{j}|X_{i},\boldsymbol{Z}))$. In practice a label-specific threshold $\theta_{j}\approx 0$ is applied to Eq (4) to identify conditional dependence:
$Y_{j}\not\!\perp X_{i}\mid\boldsymbol{Z}\quad\Leftrightarrow\quad I(Y_{j},X_{i}\mid\boldsymbol{Z})>\theta_{j}\approx 0.$ (9)
[…]
Theorem 1 (Markov Boundary Identification in Event Sequences).
If $S^{k}_{l}$ a multi-labeled sequence drawn from a dataset $D=\{S^{1}_{l},\cdots,S^{m}_{l}\}\subset\mathbb{S}$ where two Oracle Models $\text{Tf}_{x}$ and $\text{Tf}_{y}$ were trained on, then under causal sufficiency (A3), bounded lagged effects (A2) and temporal precedence (A1), the Markov Boundary of each label $Y_{j}$ in the causal graph $\mathbb{G}$ can be identified using conditional mutual information for CI-testing.
Theorem 1 enables us to sequentially recover the Markov Boundary of each label in a sequence. It provides a theoretical guarantee to recover the correct causes for each label $Y_{j}$.
Computation. A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], phase 1 is agnostic to both. Figure 2 shows all parallelised steps on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels. This transitions the time complexity from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch, with $L$ being the sequence length.
To ensure stable conditional entropy estimates and reliable predictions from $\text{Tf}_{y}$, the CMI is computed after observing $c$ events (context). This design choice also enables out-of-the-box parallelisation. By sampling $N$ variations of the prefix sequence $S_{\leq c}$, the CMI is independently computed across positions $i\in[c,L]$. In our experiments, we set $c=15,L=192$. The implementation of Phase 1 in Pytorch [pytorch] is provided in Appendix D.2.
Figure 2: The overview of CARGO. Phase 1 (One-shot) is on top, and Phase 2 (Adaptive Thresholding) is on the bottom. $d$ denotes the hidden dimension, $L$ the sequence length, $m$ the number of samples and $\text{MB}_{1},\text{MB}_{2}$ the Markov Boundary of $Y_{1},Y_{2}$. All green and blue areas are parallelised.
4 Structural Fusion of Markov Boundaries
Let $\{\mathbb{G}_{k}=(V_{k},E_{k})\}^{m}_{k=1},V_{i}\in\boldsymbol{U}$ be the set of DAGs generated by the Phase 1 from the dataset $\boldsymbol{D}$ containing $m$ i.i.d sequences $\{S^{k}_{l}\}^{m}_{k=1}$ drawn from a joint distribution $p(x,y)$. Each graph $\mathbb{G}_{k}$ represents local Markov Boundaries identified within sequence $S^{k}_{l}$. Our objective is to fuse these local graphs into a single, global consensus graph $\mathbb{G}^{*}=(\boldsymbol{U},E)$ (see Fig. 8) with the events always as parents of labels $Y_{j}$ such as:
$\text{Pa}(Y_{j})\subseteq\{X_{1},\cdots,X_{n}\}$
A naive fusion approach, such as taking the simple union of all edges $E=\bigcup^{m}_{k=1}E^{\sigma}_{k}$ works iff the Oracle models yield the perfect CI-tests in Phase 1 and thus the local graphs $\mathbb{G}^{k}$ are faithful to $p(x,y)$ ([consensus_bn_jose], Theorem 4.). Here, the ordering $\sigma$ doesn’t matter since we are dealing with Markov Boundaries. Moreover, $G^{*}$ is naturally a DAG because we considered previously that outcome labels are solely explained by events, which simplifies acyclicity. We define the Bernoulli variable $Z^{k}_{i,j}$ for each potential edge $(X_{i}\rightarrow Y_{j})$ within each sequence $S^{k}_{l}$:
$Z^{k}_{i,j}=\begin{cases}1\;\text{if the edge}\;X_{i}\rightarrow Y_{j}\;\text{is present in }\;\mathbb{G}_{k}\\ 0\;\text{otherwise}\end{cases}$
Under the Oracle Models assumption (A4), our one-shot discovery phase acts as a perfect conditional independence tester. Consequently, the detection of an edge in a local graph $\mathbb{G}_{k}$ corresponds precisely to a true causal dependency in the global graph $\mathbb{G}^{*}$. The probability of this event $P(Z^{k}_{i,j}=1)$ is therefore the true marginal probability of the edge’s existence, which we denote as $\pi_{i,j}$.
The empirical frequency, $\hat{\pi}_{i,j}(m)$, of the edge $(X_{i}\rightarrow Y_{j})$ after observing $m$ sequences is the sample mean of these i.i.d Bernoulli variables $\hat{\pi}_{i,j}(m)=\frac{1}{m}\sum_{k=1}^{m}Z_{ij}^{k}$. By the Law of Large Numbers (LLN), as the number of i.i.d sequences $m$ tends to infinity, the empirical frequency converges in probability to the true expected value of the random variable:
$\hat{\pi}_{i,j}(m)\xrightarrow{p}\mathbb{E}[Z_{i,j}^{(k)}]=\pi_{i,j}$
Thus, given a sufficiently large number of sequences, the empirical frequency $\hat{\pi}_{i,j}(m)$ serves as a consistent estimator for the true probability of the edge’s existence in the global DAG $\mathbb{G}^{*}$.
4.1 Aggregation under imperfect CI-tests
The assumption of an Oracle CI-tester, while necessary for initial theoretical guarantees, is invariably violated in practice due to factors like model capacity, limited data or class imbalance. The extracted one-shot graphs will most likely violate the independencies in $\mathbb{G}_{k}$ and thus $\mathbb{G}^{*}$.
Let us model the performance of our one-shot CI-test for any potential edge $X_{i}\rightarrow Y_{j}$ with the following error rates: (1) False Positive Rate (Type I Error): $\alpha=P(\text{detect | edge is spurious)}$ (2) True Positive Rate (Sensitivity): $1-\beta=P(\text{detect | edge is causal)}$. We operate under the reasonable assumption that our one-shot classifier is significantly better than random, which implies that $1-\beta\gg\alpha$. The expected value of our Bernoulli variable $Z_{i,j}^{k}$ is now:
$\mathbb{E}[Z_{ij}^{k}]$ $=P(Z_{ij}^{k}=1)$
$=P(\text{detect}\;|\;\text{causal})P(\text{causal})+P(\text{detect}\;|\;\text{spurious})P(\text{spurious})$
$=(1-\beta)\pi_{ij}+\alpha(1-\pi_{ij})$
The empirical frequency now converges to this new expectation. For a true edge $(\pi_{i,j}=1)$, the empirical frequency converges to a high value: $\hat{\pi}_{ij}(m)\xrightarrow{p}1-\beta$ and for spurious edge $(\pi_{i,j}=0)$ it converges to a low value: $\hat{\pi}_{ij}(m)\xrightarrow{p}\alpha$ This reveals the critical role of frequency aggregation as a mechanism for separating signal from noise.
4.2 Adaptive Fusion for Structural Discovery in Long-Tail Distributions
A primary challenge in real-world causal discovery is the long-tail distribution of outcome labels, where a few "head" labels possess abundant data while the vast majority of "tail" labels are data-sparse [longtailissue].
For rare labels, where empirical edge frequencies are high-variance estimators, a conservative high threshold is necessary to maintain precision against statistical noise. Conversely, for common labels where frequencies are reliable, a high threshold would be overly stringent, purging weaker but valid causal links. To resolve this, we introduce an adaptive thresholding strategy (Fig. 7) that tailors the edge inclusion criterion to the statistical power available for each label. We define a label-specific threshold $\tau_{j},$ as a logistic decay function of its sample support $m_{j}$:
$\tau_{j}(m_{j})=(\tau_{\max}-\tau_{\min})\cdot\frac{1}{1+e^{k(\log m_{j}-\log m_{0})}}+\tau_{\min}$ (10)
This function smoothly interpolates between a user-defined maximum threshold, $\tau_{max}$ (prioritising precision for the tail), and a minimum, $\tau_{min}$ (prioritising recall for the head). Crucially, the function’s behaviour is calibrated by the data’s distribution. Such that decay midpoint $m_{0}$ is set to the median of all label supports, providing a robust anchor point against skew.
The decay rate, $k$ is made inversely proportional to the log-inter-quartile range of supports such that $k=\frac{2\log 3}{\log q_{75}-\log q_{25}}$. For rare labels with small $m_{j}$, the high variance of the frequency estimate necessitates a high threshold $\tau_{j}(m_{j})$ that acts as a strong regularizer. For common labels with large $m_{j}$, the LLN guarantees the convergence of $\hat{\pi}_{i,j}$ to the true edge probability, justifying a lower threshold to capture a more complete causal structure. Hence, this strategy effectively functions as a data-driven denoising mechanism and shares theoretical parallels with ensembling methods [beiman_decision_trees, consistency_rate_of_dc_tree] to enhance the robustness and accuracy of the final fused graph.
Figure 3: Example of an error pattern $(y_{1})$ defined as a boolean rules of diagnosis trouble codes $(x_{i})$
$y_{1}=x_{1}\;\&\;x_{5}\;\&\;x_{8}\;\&\;x_{18}\;\&\;x_{12}\;\&\;x_{3}\;\&\;!x_{10}\;\&\;!x_{20}$
5 Empirical Evaluation
Settings & Vehicle Dataset. We used a $g4dn.12xlarge$ instance from AWS Sagemaker to run comparisons. It contains 48 vCPUs and 4 NVIDIA T4 GPUs. We used a combination of F1-Score, Precision, and Recall with different averaging [reviewmultilabellearning] to perform comparisons. We evaluated our method on a real-world vehicular test set of $m=300,000$ sequences, with $|\mathbb{Y}|=474$ different error patterns and $|\mathbb{X}|=29,100$ different DTCs forming sequences of $\approx 100\pm 35$ events. We used $\text{Tf}_{x}$ and $\text{Tf}_{y}$ with 90m and 15m parameters [math2024harnessingeventsensorydata]. The two NADEs didn’t see the test set during training. The error patterns are manually defined by domain experts as boolean rules between DTCs (Fig. 3). We set the elements of this rule as the correct Markov Boundary for each label $y_{j}$ in the tested sequences.
Multi-label Causal Discovery Comparisons. We benchmark CARGO against local structure learning (LSL) algorithms that estimate Markov Boundaries. This includes established approaches such as CMB [cmb], MB-by-MB [WANG2014252], PCD-by-PCD [pcdpcd], IAMB [iamb] from the PyCausalFS package [causality_based_feature_selection_2019], as well as the more recent, state-of-the-art MI-MCF [mimcf]. 6 random folds of the test data were created and converted into a multi-one-hot data-frame where one row represents one sequence and each column represents an event type or label $(\mathbb{X,Y})$.
Ablation on Aggregation Criterions for Phase 2. We provide an Ablation of the different criterion used in the structural fusion of Markov Boundaries. Union stands for a simple union over all edges without any removal. Frequency or edge voting, counts how often is $X_{i}\in\text{MB}(Y_{j})$. Then apply a static frequency threshold $\tau=[0.05,0.25,0.5,0.8]$. MI uses the mutual information between events and label as criterion in a Background Equivalence Search [ges]. Expected FPR (false positive ratio) [consensus_bn_freq_cutoff_no] describes two beta distributions which are fitted using the distribution of the mutual information $I(X_{i},Y_{j})$ extracted from Phase 1. The lower tail is used for outlier detection. Different FPR are chosen $\beta=[0.01,0.05,0.15,0.2]$. Detailed definitions can be found in Appendix C.4.
5.1 Results
Comparisons. We performed comparaisons on Table 1 with $n=50,000$ random sequences. We found out that even under this reduced setup, LSL algorithms failed to compute the Markov Boundaries within multiple days (3 days timeout), far exceeding practical limits for deployment. This behaviour highlights the current infeasibility of multi-label causal discovery in high-dimensional event sequences. Current algorithms are cursed under high-dimensional data since they rely on expensive CI testing that scales quadratically with the number of nodes [cd_temporaldata_review]. This positions CARGO as a more feasible approach for large-scale, multi-label causal discovery.
Table 1: Comparisons of MB retrieval with $m=50,000$ samples averaged over 66-folds with $|Y|=474,|X|=29,100$ nodes. Averaging is ’weighted’. The symbol ’-’ indicates that the algorithm didn’t output the MBs within 3 days. Metrics are given in $\%$.
Algorithm Precision Recall F1 Running Time (min)
IAMB - - - >4320>4320
CMB - - - >4320>4320
MB-by-MB - - - >4320>4320
PCDbyPCD - - - >4320>4320
MI-MCF - - - >4320>4320
CARGO $\mathbf{60.6\pm 1.5}$ $\mathbf{45.8\pm 1.7}$ $\mathbf{45.8\pm 1.2}$ $\mathbf{11.7}$
Criterions. Figure 4 illustrates the impact of aggregation choices during Phase 2. A naïve Union maximizes recall ($84\%\;\text{for weighted}$) but suffers from poor precision. When optimising a local scoring criterion as the mutual information BES mi, didn’t enhance significantly the performance over a basic Union.
Moreover, instead of optimising a score, fitting Beta distributions to detect outliers using their mutual information seems to work better; hence frequency beta is ahead, particularly with a lower FPR ratio. Frequency approaches with a static threshold confirms the analysis in Section 4.1. When a large number of samples per class $m_{j}$ is available, the frequency cut-off $\tau$ needs to be lower to not penalise classes with big support. Thus, we see that frequency with $\tau=[0.5,0.8]$ have the lower weighted f1 score of all criterions. On the otherhand, a small cut-off $\tau=[0.05,0.25]$ enables a huge improvement in the weighted average $(+40\%\;\text{precision})$, but decrease its macro average metrics $(-20\%\;\text{in precision})$.
Finally, our adaptive thresholding criterion leverages a small threshold for big supports and a big threshold for small supports, which takes advantage of the long-tail distribution. As a result, it is first on both averagings, with respectively $44.88\%\;\text{and}\;40.9\%$ for weighted and macro f1 score, and $62.8\%\;\text{and}\>66.1\%$ for weighted and macro precision.
Figure 4: Comparison of different criterions for the structural fusion (Phase 2) in function of the number of samples $m$. With $|Y|=474,|X|=29,100$ nodes.
6 Conclusion
We introduced CARGO, a novel framework for multi-label causal discovery in high-dimensional event sequences. By combining one-shot causal discovery with adaptive frequency-aware aggregation, CARGO successfully recovers interpretable causal structures from noisy observational data—achieving results in minutes where classical methods fail to scale.
CARGO could scale further by leveraging general-purpose foundation models for sequences (e.g., time-series transformers pretrained across domains). Such models could extend the applicability beyond automotive diagnostics to healthcare, cybersecurity, and other structured domains.
Under the temporal assumptions, large samples, faithfulness and perfect CI-test, CARGO recovers the true set of Markov Boundaries. However, we underlined the practical limitations, in particular long-tail distributions, a common trait in high-dimensional labeled data. Future work should extend this framework to support event-to-event causality and possibly relax assumptions such as bounded lagged effects or temporal precedence.
Ultimately, CARGO demonstrates how structured probabilistic methods can bridge the gap between causal discovery theory and scalable, practical deployment in complex industrial systems.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper aims to perform multi-labelled causal discovery from a batch of multi-labelled event sequences. It is shown in the experiment section that the current local structure learning algorithms fail to solve this causal discovery problem. In contrast, the proposed CARGO framework effectively recovers Markov Boundaries of each label, as validated on a challenging real-world dataset, aligning well with the claims made in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper addresses limitations in the Conclusion section. Notably, Assumption A4 regarding Oracle models can be sensitive to model capacity and training data, as shown in the Table 2. Additionally, while the evaluation relies on a single high-dimensional multi-labelled event sequence dataset, it is reflective of real-world complexity (e.g., event sparsity and high cardinality), which is sufficient to prove that current methods are not well adapted. Moreover, the absence of publicly available datasets with clearly defined Markov Boundaries limits for high-dimensional data makes the direct comparison limited. Still, CARGO’s ability to identify meaningful causal structures without strict ground-truth labels remains a significant advantage in practice.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Under the shown assumptions, the Lemma 1 and 2 proofs are given in the Appendix B.1, B.2. Moreover, proof of Theorem 1 is given in Appendix B.3 and the frequency section provides a guarantee under a large number of samples for each class.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The general settings and configuration are provided in Section 5. The code for CARGO are provided anonymously in the Appendix D.2 and D.3. Moreover, alongside the code, the dataset used for evaluation will be provided upon acceptance, which is the first large-scale high-dimensional dataset about modern vehicle diagnostics leading to error patterns that is made public. The two Transformers models are described in [math2024harnessingeventsensorydata].
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in the supplemental material?
  23. Answer: [Yes]
  24. Justification: A code example is provided in the Appendix D.2 and D.3. Moreover, alongside the code, the dataset used for evaluation will be provided upon acceptance, which is the first large-scale high-dimensional dataset about modern vehicle diagnostics, leading to error patterns that is made public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The data characteristics, such as the number of sequences, events, labels, mean and standard deviation of the sequence length, are provided in the experiment section. Hyperparameters such as the sampling number $N$ or the confidence parameter $k$ are provided as ablation studies in Appendix C.2, C.3, as well as the NADEs configuration in Table 2.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Results of the experiments are reported using mean and one standard deviation after 6 runs on 6 random data samples (Table 1). The classification performance as a function of the number of samples is reported using 1-sigma error bars in Figure 4 and shown for two different averaging methods across several runs on the same data sample. Ablation studies are also given using 1-sigma error bars to appropriately choose the hyperparameters in Section C.2, C.3.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The exact computational resources are provided in Section 5 as well as the running time of the different experiments in Tables 1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This research does not involve human participants.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: As pointed out in the introduction, related work and conclusion, impacts should be in decision-making areas like diagnostics, manufacturing, healthcare or cybersecurity. If the assumption were carefully evaluated, CARGO would enhance decision-making and explainability of causal relationships between events and outcome labels.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There are no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: The paper and method for both phases are currently patented.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The CARGO code, models, and pre-trained assets are released under the Creative Commons Attribution-NonCommercial 4.0 International License (CC BY-NC 4.0) License, with ongoing patent protection for commercial use. The full documentation is provided alongside the code to ensure reproducibility and transparency, including model training details, dataset preparation, and usage instructions. Unauthorised commercial use, reproduction, or distribution of this work without the explicit written consent of the patent holder is strictly prohibited.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The associated experiments are based on anonymised datasets. No crowdsourcing or direct human subject involvement was required to collect or validate the data. As such, no direct participant interactions or compensations were involved.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve human subject research.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLMs were used for causal discovery as explained in the methodology section.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Notations and Definitions
A.1 Notations
We use capital letters (e.g., $X$) to denote random variables, lower-case letters (e.g., $x$) for their realisations, and bold capital letters (e.g., $\boldsymbol{X}$) for sets of variables. Let $\boldsymbol{U}$ denote the set of all (discrete) random variables. We define the event set $\boldsymbol{X}=\{X_{1},\ldots,X_{n}\}\subset\boldsymbol{U}$, and the label set $\boldsymbol{Y}=\{Y_{1},\ldots,Y_{n}\}\subset\boldsymbol{U}$. When explicitly said, event $X^{(t_{i})}_{i}$ represent the occurrence of $X_{i}$ at the sequence step $i$ and time $t_{i}$. Similarly for $Y^{(t_{i+1})}_{i+1}$.
A.2 Definitions
Definition 1 (Faithfulness).
Spirtes2001CausationPA. Given a BN $\boldsymbol{U},\mathbb{G},P>,\mathbb{G}$ is faithful to $P$ if and only if every conditional independence present in $P$ is entailed by $\mathbb{G}$ and the Markov condition holds. $P$ is faithful if and only if there exist a DAG $\mathbb{G}$ such that $\mathbb{G}$ is faithful to $P$.
Definition 2 (Conditional Independence).
Variables $X$ and $Y$ are said to be conditionally independent given a variable set $\boldsymbol{Z}$, if $P(X,Y|\boldsymbol{Z})=P(X|\boldsymbol{Z})P(Y|\boldsymbol{Z})$, denoted as $X\bot Y|\boldsymbol{Z}$. Inversely, $X\not\perp Y|\boldsymbol{Z}$ denotes the conditional dependence. Using the conditional mutual information [cover1999elements] to measure the independence relationship, this implies that $\text{I}(X,Y|\boldsymbol{Z})=0\Leftrightarrow X\perp Y|\boldsymbol{Z}$.
Definition 3 (Markov Boundary).
optimal_feature_set_cd. In a faithful BN $\boldsymbol{U},\mathbb{G},P>$, for a set of variables $\boldsymbol{Z}\subset\boldsymbol{U}$ and label $Y\in\boldsymbol{U}$, if all other variables $X\in\{\boldsymbol{X}-\boldsymbol{Z}\}$ are independent of $Y$ conditioned on $\boldsymbol{Z}$, and any proper subset of $\boldsymbol{Z}$ do not satisfy the condition, then $\boldsymbol{Z}$ is the Markov Boundary of $Y$: $\textbf{MB}(Y)$.
Definition 4.
(Markov Equivalence Class). Two distinct graph $\mathbb{G},\mathbb{G^{\prime}}$ are said to belong to the same Markov Equivalence Class (MEC) if they have the same set of conditional independencies i.e $I(\mathbb{G})=I(\mathbb{G^{\prime}})$.
Definition 5.
(Decomposable Criterion). We say that a scoring criterion $S(\mathbb{G},D)$ is decomposable if it can be written as a sum of measures, each of which is a function only of one node and its parents. In other words, a decomposable scoring criterion $\mathbb{S}$ applied to a DAG $\mathbb{G}$ can always be expressed as:
$S(\mathbb{G},\boldsymbol{D})=\sum^{n}_{i}s(X_{i},\textbf{Pa}^{\mathbb{G}}_{i})$ (11)
Definition 6 (Score equivalent).
ges. A score $S$ is score equivalent if it assigns the same score to all the graphs in the same MEC.
Definition 7 (Local Consistency).
ges Let $\boldsymbol{D}$ contain $m$ iid samples from some distribution $p(.)$. Let $\mathbb{G}$ be any possible DAG and $\mathbb{G}^{\prime}$ a different DAG obtained by adding the edge $i\rightarrow j$ to $\mathbb{G}$. A score $S$ is locally consistent if both hold:
  • If $X_{i}\not\perp_{p}X_{j}|Pa^{\mathbb{G}}_{j},$ then $S(\mathbb{G^{\prime}},\boldsymbol{D})>S(\mathbb{G},\boldsymbol{D})$
  • If $X_{i}\perp_{p}X_{j}|Pa^{\mathbb{G}}_{j},$ then $S(\mathbb{G^{\prime}},\boldsymbol{D})<S(\mathbb{G},\boldsymbol{D})$
A.3 Assumptions
Assumption 1 (Temporal Precedence).
Given a perfectly recorded sequence of events $((x_{1},t_{1}),\cdots,(x_{L},t_{L}))$ with labels $(\boldsymbol{y}_{L},t_{L})$ and monotonically increasing time of occurrence $0\leq t_{1}\leq\cdots\leq t_{L}$, an event $x_{i}$ is allowed to influence any subsequent event $x_{j}$ such that $t_{i}\leq t_{j}$ and $i<j$. Formally, the graph $\mathbb{G}=(\boldsymbol{U},\boldsymbol{E})$, $(x_{i},x_{j})\in\boldsymbol{E}\implies t_{i}\leq t_{j}\;\text{and step}\;i<j$
It allows us to remove ambiguity in causal directionality, and is a widely used assumption across time-series and sequential data [cd_temporaldata_review].
Assumption 2 (Bounded Lagged Effects).
Once we observed events up to timestamp $t_{i}$ and step $i$ as $\boldsymbol{Z}_{\leq t_{i}}=((x_{1},t_{1}),\cdots,(x_{i},t_{i}))$, any future lagged copy of event $X^{(t_{i}+\tau)}_{i}$ is independent of $Y_{j}$ conditioned on $\boldsymbol{Z}_{\leq t_{i}}$:
$Y_{j}\perp X^{(t_{i}+\tau)}_{i}|\boldsymbol{Z}_{\leq t_{i}}$
Where $\tau=t_{i+1}-t_{i}$ is a finite bound on the allowed time delay for causal influence.
In other words, we allow the causal influence of event $X_{i}$ on $Y_{j}$ until the next event $X_{i+1}$ is observed. We note that for data with strong lagged effects (e.g., financial transactions), this might not hold well, but for log-based and error code-based data this is usually correct.
Assumption 3 (Causal Sufficiency for Labels).
All relevant variables are observed, and there are no hidden confounders affecting the labels.
Assumption 4 (Oracle Models).
We assume that two autoregressive Transformer models, $\text{Tf}_{x}$ and $\text{Tf}_{y}$, are trained via maximum likelihood on a dataset of multi-labeled event sequences $D=\{S^{1}_{l},\cdots,S^{m}_{l}\}\subset\mathbb{S}$, and can perfectly approximate the true conditional distributions of events and labels:
$P(X_{i}|\text{Pa}(X_{i}))=P_{\theta_{x}}(X_{i}|\text{Pa}(X_{i}))=\text{Tf}_{x}(S_{<i}),\;P(Y_{j}|\text{Pa}(Y_{j}))=P_{\theta_{y}}(Y_{j}|\text{Pa}(Y_{j}))=\text{Tf}_{y}(S_{\leq i})$ (12)
A.4 Lemmas
Lemma 1 (Identifiability of $\mathbb{G}$).
Assuming the faithfulness condition holds for the true causal graph $\mathbb{G}$. Let $\text{Tf}_{x}$ and $\text{Tf}_{y}$ be oracle models that model the true conditional distributions of events and labels, respectively. The joint distribution $P_{\theta_{x},\theta_{y}}$ can then be constructed, and any conditional independence detected from the distributions estimated by $\text{Tf}_{x}$ and $\text{Tf}_{y}$ corresponds to a conditional independence in $\mathbb{G}$:
$X_{i}\perp_{\theta_{x},\theta_{y}}Y_{j}\mid\boldsymbol{Z}\quad\implies\quad X_{i}\perp_{\mathbb{G}}Y_{j}\mid\boldsymbol{Z}.$
Where $\perp_{\theta_{x},\theta_{y}}$ denotes the independence entailed by the joint probability $P_{\theta_{x},\theta_{y}}$.
Lemma 2 (Markov Boundary Equivalence).
In a multi-label event sequence $S_{l}$ and under the temporal precedence assumption A1, the Markov Boundary of each label $Y_{j}$ is only its parents such that $\forall X\in\{\boldsymbol{U}-\text{Pa}(Y_{j})\},X\perp Y_{j}|\text{Pa}(Y_{j})\Leftrightarrow\text{MB}(Y_{j})=\text{Pa}(Y_{j})$.
Appendix B Proofs
We provide proofs for the results described in Section 3
B.1 Proof of Lemma 1
Proof.
We assume that the data is generated by the associated causal graph $\mathbb{G}$ following the sequential BN from a multi-labelled sequence $S$. And that the faithfulness assumption holds [pearl_1998_bn], meaning that all conditional independencies in the observational data are implied by the true causal graph $\mathbb{G}$.
Given that the Oracle models $\text{Tf}_{x}$ and $\text{Tf}_{y}$ are trained to perfectly approximate the true conditional distributions, for any variable $U_{i}$ in the graph, we have:
$P(U_{i}|\text{Pa}(U_{i}))=\begin{cases}P(Y_{j}|\text{Pa}(Y_{j}))=P_{\theta_{y}}(Y_{j}|\text{Pa}(Y_{j})),&\text{if }U_{i}\in\boldsymbol{Y}\\ P(X_{i}|\text{Pa}(X_{i}))=P_{\theta_{x}}(X_{i}|\text{Pa}(X_{i})),&\text{otherwise}.\end{cases}$
The joint distribution $P_{\theta_{x},\theta_{y}}$ can then be constructed using the chain rule $P_{\theta_{x},\theta_{y}}(X_{1},\cdots,X_{i},Y_{1},\cdots,Y_{c})=\prod^{i}_{k=0}P(X_{k}|Pa(X_{k}))\prod^{c}_{l}P(Y_{l}|\text{Pa}(Y_{l})$. By the faithfulness assumption [pearl_1998_bn], if the conditional independencies hold in the data, they must also hold in the causal graph $\mathbb{G}$:
$X_{i}\perp Y_{j}|\boldsymbol{Z}\implies X_{i}\perp_{\mathbb{G}}Y_{j}|\boldsymbol{Z}$
Since we can approximate the true conditional distributions, it follows that:
$X_{i}\perp_{\theta_{x},\theta_{y}}Y_{j}|\boldsymbol{Z}\implies X_{i}\perp Y_{j}|\boldsymbol{Z}\implies X_{i}\perp_{\mathbb{G}}Y_{j}|\boldsymbol{Z}$
Where $\perp_{\theta_{x},\theta_{y}}$ denotes the independence entailed by the joint probability $P_{\theta_{x},\theta_{y}}$. Thus, the graph $\mathbb{G}$ can be identified from the observational data. ∎
B.2 Proof of Lemma 2
Proof.
Let $\boldsymbol{U},\mathbb{G},P>$ be the sequential BN composed of the events from the multi-labeled sequence $S_{l}=(\{(t_{1},x_{1},\cdots,(t_{L},x_{L})\}^{L}_{i=1},(\boldsymbol{y}_{L},t_{L}))$. Following the temporal precedence assumption A1, the labels $\boldsymbol{y_{L}}$ can only be caused by past events $(x_{1},\cdots,x_{L})$; moreover by definition, labels do not cause any other labels. Thus, $Y_{j}$ has no descendants, so no children and spouses. Therefore, together with the Markov Assumption we know that $\forall X\in\{\boldsymbol{U}-Pa(Y_{j})\}:Y_{j}\perp X|Pa(Y_{j})$. Which is the definition of the MB (Def. 3). Thus, $\textbf{MB}(Y_{j})=Pa(Y_{j})$.
B.3 Proof of Theorem 1.
Proof.
By recurrence over the sequence length $L$ of the multi-label sequence $S^{k}_{l}$, we want to show that under temporal precedence A1, bounded lagged effects A2, causal sufficiency A3, Oracle Models A4 the Markov Boundary of label $Y_{j}$ can be identified in the causal graph $\mathbb{G}$.
Let’s define $\mathcal{M}^{L}_{j}$ as the estimated Markov Boundary of $Y_{j}$ after observing $L$ events.
Base Case: $\mathbf{L=1}$: Consider the BN for step $L=1$ following the Markov assumption [pearl_1998_bn] with two nodes $X_{1},Y_{j}$. Using $\text{Tf}_{x},\text{Tf}_{y}$ as Oracle Models A4, we can express the conditional probabilities for any node $U$:
$P(U|\text{Pa}(U))=\begin{cases}P(X_{1})=P_{\theta_{x}}(X_{1}|[CLS])\;\text{if}\;U\in\boldsymbol{X}\\ P(Y_{j}|X_{1})=P_{\theta_{y}}(Y_{j}|X_{1})\;\text{otherwise}\end{cases}$ (13)
Assuming that P is faithful (A1) to $\mathbb{G}$, no hidden confounders bias the estimate (A3) and temporal precedence (A1), we can estimate the CMI 4 such that iif $I(X_{1},Y_{j})|\emptyset)>0\Leftrightarrow Y_{j}\not\perp_{\theta_{x},\theta_{y}}X_{1}\implies Y_{j}\not\perp_{\mathbb{G}}X_{1}$ (Lemma 1).
Since we assume temporal precedence A1, we can orient the edge such that $X_{1}$ must be a parent of $Y_{j}$ in $\mathbb{G}$. Using Lemma 2, we know that $Par(Y_{j})=\textbf{MB}(Y_{j})\implies X_{1}\in\textbf{MB}(Y_{j})$, thus we must include $X_{1}$ in $M^{1}_{j}$, otherwise not.
Heredity: For $L=i$, we obtained $M^{i}_{j}$ with the sequential BN up to step $L=i$. Now for $L=i+1$, the sequential BN has $i+2$ nodes denoted as $\boldsymbol{U^{\prime}}=(X_{1},\cdots,X_{i},X_{i+1},Y_{j})$. Using the Oracle Models A4 and following the Markov assumption [pearl_1998_bn], we can estimates the following conditional probabilities for any nodes $U\in\boldsymbol{U^{\prime}}$:
$P(U|\text{Pa}(U))=\begin{cases}P(Y_{j}|\text{Pa}(Y_{j}))\approx P_{\theta_{y}}(Y_{j}|\text{Pa}(Y_{j})),&\text{if }U\in\boldsymbol{Y}\\ P(X|\text{Pa}(X))\approx P_{\theta_{x}}(X|\text{Pa}(X)),&\text{otherwise}.\end{cases}$ (14)
By bounded lagged effects (A2) we know that the causal influence of past $X_{\leq i}$ on $Y_{j}$ has expired. Moreover, no hidden confounders (A3) bias the independence testing. Finally, using Eq. (4) we can estimate the CMI such that iif $I(Y_{j},X_{i+1}|\boldsymbol{Z})>0\Leftrightarrow Y_{j}\not\perp_{\theta_{x},\theta_{y}}X_{i+1}|\boldsymbol{Z}\implies Y_{j}\not\perp_{\mathbb{G}}X_{i+1}|\boldsymbol{Z}$ (Lemma 1).
Since we assume temporal precedence A1, we can orient the edge so that $X_{i+1}$ must be a parent of $Y_{j}$ in $\mathbb{G}$. Using Lemma 2, we know that $Par(Y_{j})=\textbf{MB}(Y_{j})\implies X_{i+1}\in\textbf{MB}(Y_{j})$. Thus $X_{i+1}\in M^{i+1}_{j}$ which represent the MB($Y_{j})$ for step $i+1$.
Finally, $\mathcal{M}^{i+1}_{j}$ still recovers the Markov Boundary of $Y_{j}$ such that
$\forall U\in\{\boldsymbol{U^{\prime}}-\mathcal{M}^{i+1}_{j}\},Y_{j}\perp U|\mathcal{M}^{i+1}_{j}$
Appendix C Ablations
C.1 NADEs Quality.
We did several ablations on the quality of the NADEs and their impact on the one-shot causal discovery phase. In particular, Table 2 presents multiple $\text{Tf}_{x},\text{Tf}_{y}$ with respectively 90 and 15 million parameters or 34 and 4 million parameters. We also varied the context window (conditioning set $\boldsymbol{Z}$), trained on different amounts of data (tokens) and reported the classification results on the test set of $\text{Tf}_{y}$ alone. We didn’t output the running time since it was approximately the same for all NADEs: $1.27$ minutes of 50,000 samples and $0.14$ for 5000.
We observe that scaling up the NADEs model size and the trained data show the biggest improvements. After, it is via the context $c$, which after $c=15$ shows declined performance for a larger number of samples. We then choose the backbone with 1.5B Tokens, 105m parameters and a context $c=15$ for our experiments.
Table 2: Ablations of the performance of Phase 1 (One-shot MB retrieval) in function of different NADEs with m=50,000m=50{,}000 and $m=500$ samples averaged over 6-folds. Classification metrics use weighted averaging. Metrics are given in $\%$.
Tokens Parameters Context Precision (↑) Recall (↑) F1 Score (↑) Tfy F1 (↑)
For n=50,000n=50{,}000 samples
1.5B 105m $c=4$ $47.95\pm 1.05$ $30.65\pm 0.51$ $37.39\pm 0.67$ 88.6
1.5B 105m $c=12$ $54.62\pm 1.03$ $29.88\pm 0.73$ $38.63\pm 0.85$ 90.43
1.5B 105m $c=15$ $\mathbf{55.26\pm 1.42}$ $\mathbf{31.37\pm 0.82}$ $\mathbf{40.02\pm 1.03}$ 90.57
1.5B 105m $c=20$ $49.52\pm 1.59$ $\mathbf{31.76\pm 0.85}$ $36.54\pm 1.10$ 91.19
1.5B 105m $c=30$ $36.65\pm 1.18$ $22.75\pm 0.78$ $26.57\pm 0.91$ 92.64
300m 47m $c=20$ $39.49\pm 1.77$ $26.30\pm 0.89$ $29.01\pm 1.10$ 83.6
For $n=500$ samples
1.5B 105m $c=12$ $54.84\pm 4.55$ $\mathbf{31.45\pm 2.23}$ $\mathbf{39.95\pm 2.83}$ 90.43
1.5B 105m $c=15$ $55.04\pm 3.36$ $29.90\pm 1.78$ $38.74\pm 2.24$ 90.57
1.5B 105m $c=20$ $48.84\pm 4.01$ $\mathbf{31.65\pm 2.37}$ $36.19\pm 2.65$ 91.19
300m 47m $c=20$ $38.23\pm 2.91$ $25.31\pm 2.39$ $27.92\pm 2.25$ 83.6
C.2 Sampling Number
We experimented with different numbers of $n$ for the sampling method across different averaging (micro, macro, weighted), Fig. 5. We performed 8 different runs and reported the average, standard deviation and elapsed time. We can say that generally, sampling with a bigger $N$ tends to decrease the standard deviation and give more reliable Markov Boundary estimation. Moreover, as we process more samples, the model is gradually improving at a logarithmic growth until it converges to a final score. We also verify that our time complexity is linear with the number of samples $N$. Based on these results, we choose generally $N=68$ as the number of samples.
Figure 5: Evolution of several classification metrics (one-shot) and elapsed time per sample in function of the number of samples NN chosen. Results are reported using 1-sigma error bar.
C.3 Dynamic Thresholding
We performed ablations on the effect of $k$ during the dynamic thresholding of the CMI (Eq. (9)) to access conditional independence in Fig. 6. To balance the classification metrics across the different averaging, we set $k=2.75$.
Figure 6: Evolution of one-shot F1 Score, Precision and Recall in function of coefficient $k$. Results are reported using 1-sigma error bar.
C.4 Criterions
This section presents the different criteria used for comparison in the experimental evaluation.
C.5 Frequency
Frequency-based heuristics that apply fixed thresholds $\tau$ to the empirical frequency of the occurrence of $X_{i}$ in each of the Markov Boundary $\textbf{MB}(Y_{j})$. Formally, For each label $Y_{j}$, after merging all local edge sets into a global set $E=\bigcup^{m}_{i=1}E_{i}$, we evaluate each candidate variable $X_{i}\in\textbf{MB}_{j}$ based on its frequency of appearance across the local models. If this frequency exceeds the threshold $\tau$, the variable is retained in the final merged $\textbf{MB}_{j}$; otherwise, it is discarded.
C.5.1 Expected FPR Adjustment
Same principle as in [consensus_bn_freq_cutoff_no] except that we fit the two Beta distributions on the mutual information of $I(Y_{j},X_{i})$ instead of the raw frequencies.
C.5.2 Mutual Information
BES
The BES is the second phase of GES [ges], where edges are removed one after the other to maximise a criterion $S$. Heuristics approaches [approx_fusionPUERTA2021155, delSagrado2001] aim to solve this problem by optimizing:
$E=\underset{E_{i}\in\varepsilon}{\arg}{\text{max}}\sum_{e\in E_{i}}S(e)$ (15)
Where $\varepsilon$ denotes the search space (all possible edges over $\boldsymbol{U}$) and $S(e)$ a criterion function for edge relevance (e.g. edge frequency, thresholds, $\cdots$). This formulation takes into account the underlying edges’ characteristics but not the overall network structure, complexity and missing data [consensus_bn_jose], leading to a consensus fusion approach [torrijos2025informedgreedyalgorithmscalable].
Estimating Mutual Information in Event Sequences.
We would like to reuse the estimated conditional mutual information, Eq. (4) and profit from the parallelised inference of Phase 1 (Fig. 2).
As argued out by JanzingDominik2013QCI, a causal strength measure (or criterion) $C_{X_{i}\rightarrow Y_{j}}$ should possess multiple properties. Notably that if $C_{X_{i}\rightarrow Y_{j}}=0$, then the joint distribution satisfies the Markov condition with respect to the DAG obtained by removing the arrow $X_{i}\rightarrow Y_{j}$}. Moreover, the true DAG reads $X_{i}\rightarrow Y$ iff $C_{X_{i}\rightarrow Y_{j}}=I(X_{i},Y_{j})$.
It comes as a natural criterion to merge edges from multiple causal graphs. However, it remains tricky to estimate [estimating_mi]. Using the chain rule of conditional mutual information [cover1999elements], we can rewrite it as:
$I(Y_{j},X_{i}|\boldsymbol{Z})=I(Y_{j},X_{i})-I(Y_{j},X_{i},\boldsymbol{Z})$ (16)
Where $I(Y_{j},X_{i},\boldsymbol{Z})$ is the interaction information [cover2006elements], which tells us wether knowing $\boldsymbol{Z}$ explains away the dependency between $X_{i}$ and $Y_{j}$ (negative interaction), or enhances it (positive interaction):
$I(Y_{j},X_{i},\boldsymbol{Z})\triangleq I(Y_{j},\boldsymbol{Z})-I(Y_{j},\boldsymbol{Z}|X_{i})$
$I(Y_{j},\boldsymbol{Z})$ can be estimated using the same Monte-Carlo sampling as for $I(Y_{j},X_{i}|\boldsymbol{Z})$ (4). Since $I(Y_{j},\boldsymbol{Z})=H(Y_{j})-H(Y|\boldsymbol{Z_{j}})$, the marginal $p(y)$ is needed. Fortunately, the dataset $\boldsymbol{D}$ is large enough, hence the frequencies of $y_{j}$ are recovered empirically and an estimate $\hat{p}(y)$ which we assume to be equal to the true marginal $p(y)$. We acknowledge that under a restricted dataset, $\hat{p}(y)$ might differ from $p(y)$. This yields to:
$I(Y_{j},\boldsymbol{Z})=\mathbb{E}_{z}D_{KL}(P(Y_{j}|\boldsymbol{Z})||\hat{P}(Y_{j}))=\mathbb{E}_{z}I_{G}(Y_{j},z)$ (17)
Formally, we assume that for long sequences i.e $i\rightarrow+\infty$, our event sequences form a stationary ergodic stochastic process and $I(Y_{j},\boldsymbol{Z}|X_{i})$ is negligible compare to $I(Y_{j},\boldsymbol{Z})$ since $\boldsymbol{Z}$ is containing most of the information to predict $Y_{j}$. This reduces the mutual information to
$I(Y_{j},X_{i})\approx I(Y_{j},X_{i}|\boldsymbol{Z})+I(Y_{j}|\boldsymbol{Z})$
Criterion.
We propose a Class-Aware Information Gain (CAIG) score for evaluating candidate edges during Phase 2 of CARGO. Given $m$ i.i.d. samples from a dataset $\boldsymbol{D}$, CAIG balances three key factors: mutual information derived from information gain, class imbalance, and network complexity.
For each label node $Y_{j}$, with candidate parent set $\textbf{Pa}^{\mathbb{G}}_{j}$, the CAIG score is:
$S(\mathbb{G}^{\prime},\boldsymbol{D})=\sum_{j=1}^{n}s_{I}(Y_{j},\textbf{Pa}^{\mathbb{G}}_{j})-\alpha\cdot|\textbf{Pa}^{\mathbb{G}}_{j}|\cdot\log{\left(\frac{m}{m_{j}}+1\right)}$ (18)
With $s_{I}(Y_{j},\textbf{Pa}^{\mathbb{G}}_{j})=\sum_{X_{i}\in\textbf{Pa}^{\mathbb{G}}_{j}}I(Y_{j},X_{i})$, $\alpha$ is a regularization hyperparameter, $m_{j}$ is the number of positive instances for class $Y_{j}$.
This formulation encourages informative yet parsimonious graph structures, correcting for underrepresented labels via the regularisation term. It is also efficient since CAIG is decomposable [ges] like BIC with the local $s_{I}$. This criterion is denoted as BES mi imbalance par in our experiments in Fig. 4.
Appendix D Implementation
D.1 Computation.
A key advantage of our approach is its scalability. Unlike traditional methods whose complexity depends on the event and label cardinality $|\mathbb{X}|$ and $|\mathbb{Y}|$ [feature_selection_review], our method is agnostic to both. As illustrated in Figure 2, all steps are parallelised on GPUs. CMI estimations are independently performed for all positions $i\in[c,L]$, with the sampling pushed into the batch dimension and results averaged across labels, leading to $\text{BS}\times N\times L$ CI-tests per batch $D=\{S^{0}_{l},\ldots,S^{m}_{l}\}$. Consequently, time complexity transitions from $\mathcal{O}(\text{BS}\times N\times L)$ to $\mathcal{O}(1)$ per batch due to GPU parallelism. The complexity is bounded by the Transformers’ inference part, where it scales quadratically with the sequence length $\mathcal{O}(L^{2})$ if one uses vanilla self-attention [tf].
D.2 Phase 1
The following is the implementation of the one-shot phase in PyTorch [pytorch].
1def topk_p_sampling(z, prob_x, c: int, n: int = 64, p: float = 0.8, k: int = 35,
2 cls_token_id: int = 1, temp: float = None):
3 # Sample just the context
4 input_ = prob_x[:, :c]
5
6 # Top-k first
7 topk_values, topk_indices = torch.topk(input_, k=k, dim=-1)
8
9 # Top-p over top-k values
10 sorted_probs, sorted_idx = torch.sort(topk_values, descending=True, dim=-1)
11 cum_probs = torch.cumsum(sorted_probs, dim=-1)
12 mask = cum_probs > p
13
14 # Ensure at least one token is kept
15 mask[..., 0] = 0
16
17 # Mask and normalize
18 filtered_probs = sorted_probs.masked_fill(mask, 0.0)
19 filtered_probs += 1e-8 # for numerical stability
20 filtered_probs /= filtered_probs.sum(dim=-1, keepdim=True)
21
22 # Unscramble to match the original top-k indices
23 # Need to reorder the sorted indices back to the original top-k
24 reorder_idx = torch.argsort(sorted_idx, dim=-1)
25 filtered_probs = torch.gather(filtered_probs, -1, reorder_idx)
26
27 batched_probs = filtered_probs.unsqueeze(1).repeat(1, n, 1, 1) # (bs, n, seq_len, k)
28 batched_indices = topk_indices.unsqueeze(1).repeat(1, n, 1, 1) # (bs, n, seq_len, k)
29
30 sampled_idx = torch.multinomial(batched_probs.view(-1, k), 1) # (bs*n*seq_len, 1)
31 sampled_idx = sampled_idx.view(-1, n, c).unsqueeze(-1)
32
33 sampled_tokens = torch.gather(batched_indices, -1, sampled_idx).squeeze(-1)
34 sampled_tokens[..., 0] = cls_token_id
35
36 # Reconstruct full sequence
37 z_expanded = z.unsqueeze(1).repeat(1, n, 1)[..., c:]
38 return torch.cat((sampled_tokens, z_expanded), dim=-1)
39
40from torch import $n$
41def OneShotCD(tfe: $n$.Module, tfy: $n$.Module, batch: dict[str, torch.Tensor], c: int, n: int, eps: float=1e-6, topk: int=20, k: int=2.75, p=0.8) -> torch.Tensor:
42 """ tfe, tfy: are the two autoregressive transformers (event type and label)
43 batch: dictionary containing a batch of input_ids and attention_mask of shape (bs, L) to explain.
44 c: scalar number defining the minimum context to start inferring, also the sampling interval.
45 n: scalar number representing the number of samples for the sampling method.
46 eps: float for numerical stability
47 topk: The number of top-k most probable tokens to keep for sampling
48 k: Number of standard deviations to add to the mean for dynamic threshold calculation
49 p: Probability mass for top-p nucleus
50 """
51 o = tfe(attention_mask=batch[’attention_mask’], input_ids=batch[’input_ids’])[’prediction_logits’] # Infer the next event type
52 x_hat = torch.nn.functional.softmax(o, dim=-1)
53
54 b_sampled = topk_p_sampling(batch[’input_ids’], x_hat, c, k=topk, n=n, p=p) # Sampling up to (bs, n, L)
55 n_att_mask = batch[’attention_mask’].unsqueeze(1).repeat(1, n, 1)
56
57 with torch.inference_mode():
58 o = tfy(attention_mask=n_att_mask.reshape(-1, b_sampled.size(-1)), input_ids=b_sampled.reshape(-1, b_sampled.size(-1))) # flatten and infer
59 prob_y_sampled = o[’ep_prediction’].reshape(b_sampled.size(0), n, batch[’input_ids’].size(-1)-c, -1) # reshape to (bs, n, L-c)
60
61 # Ensure probs are within (eps, 1-eps)
62 prob_y_sampled = torch.clamp(prob_y_sampled, eps, 1 - eps)
63
64 y_hat_i = prob_y_sampled[..., :-1, :] # P(Yj|z)
65 y_hat_iplus1 = prob_y_sampled[..., 1:, :] # P(Yj|z, x_i)
66
67 # Compute the CMI & CS and average across sampling dim
68 cmi = torch.mean(y_hat_iplus1*torch.log(y_hat_iplus1/y_hat_i)+ (1-y_hat_iplus1)*torch.log((1-y_hat_iplus1)/(1-y_hat_i)), dim=1)
69 # (BS, L, Y)
70 cs = y_hat_iplus1 - y_hat_i
71 cs_mean = torch.mean(cs, dim=1)
72 cs_std = torch.std(cs, dim=1)
73
74 # Confidence interval for threshold
75 mu = cmi.mean(dim=1)
76 std = cmi.std(dim=1)
77 dynamic_thresholds = mu + std * k
78
79 # Broadcast to select an individual dynamic threshold
80 cmi_mask = cmi >= dynamic_thresholds.unsqueeze(1)
81
82 cause_token_indices = cmi_mask.nonzero(as_tuple=False)
83 # (num_causes, 3) --> each row is [batch_idx, position_idx, label_idx]
84 return cause_token_indices, cs_mean, cs_std, cmi_mask
Remark.
Since tfy contains tfe as backbone, in practice we need only one forward pass from tfy and extract also $\hat{\boldsymbol{x}}$, so tfe is not needed. We let it to improve understanding and clarity.
D.3 Phase 2.
1import random
2def create_auto_adaptive_threshold_fn(all_m_j, tau_max=0.5, tau_min=0.05, k=None, m0="median"):
3 m_0 = np.median(all_m_j)
4
5 if k == None:
6 q25, q75 = np.percentile(all_m_j, [25, 75])
7 if q75 == q25:
8 k = 1.0
9 else:
10 log_iqr = np.log(q75) - np.log(q25)
11 k = (2 * np.log(3)) / log_iqr
12
13 def threshold_function(m_j):
14 log_m_j = np.log(m_j + 1e-9)
15 log_m_0 = np.log(m_0)
16 logistic_decay = 1 / (1 + np.exp(k * (log_m_j - log_m_0)))
17 return (tau_max - tau_min) * logistic_decay + tau_min
18
19 return threshold_function
20
21def adaptive_thresholding_frequency(graphs: list,
22 present_labels: dict,
23 frequency_threshold: float = 0.5,
24 k: float=None,
25 tau_min: float=0.05,
26 tau_max: float=0.5,
27 m0: str="median",
28 verbose=False,
29 **kwargs):
30 """
31 Frequency voting: keep edges appearing with frequency > threshold across samples.
32
33 :param graphs: list of local graphs (e.g., from Phase 1). Each graph is a dict[label][token] = list of stats.
34 :param present_labels: labels present in evaluation
35 :param frequency_threshold: e.g. 0.5 for majority, 0.8 for conservative
36 :return: filtered_labels, sample_per_label, elapsed_time
37 """
38 start_time = datetime.now()
39 # Step 1: Aggregate graphs
40 labels, sample_per_label = union(graphs) # user-defined union function
41 old_labels = labels.copy()
42 nodes = count_nodes(labels)
43 samples = len(graphs)
44
45 # Create threshold function
46 auto_threshold_fn = create_auto_adaptive_threshold_fn(list(sample_per_label.values()), k=k, tau_max=tau_max, tau_min=tau_min, m0=m0)
47
48 # Step 2: Frequency voting with dynamic thresholds
49 edge_counts = defaultdict(lambda: defaultdict(int)) # edge_counts[label][token] = count
50
51 for g in graphs:
52 for label, token_dict in g.items():
53 if label not in labels:
54 continue
55 for token in token_dict:
56 edge_counts[label][token] += 1
57
58 # Step 3: Keep edges above frequency threshold
59 filtered_labels = defaultdict(dict)
60 for label in labels:
61 total = sample_per_label.get(label, samples) # fallback to total graphs if missing
62 for token, count in edge_counts[label].items():
63 freq = count / total
64 if freq >= auto_threshold_fn(sample_per_label.get(label, 1)):
65 filtered_labels[label][token] = {’frequency’: freq}
66 if verbose:
67 print(f"[{label}] token {token} kept (freq={freq:.2f})")
68
69 nb_of_edges = sum(len(v) for v in filtered_labels.values())
70 print(f"Time: {(datetime.now() - start_time).total_seconds():.2f}s")
71 return filtered_labels, sample_per_label, (datetime.now() - start_time).total_seconds()
Appendix E Figures
Figure 7: Adaptive thresholding function $\tau_{j}(m_{j})$ across varying label frequencies $m_{j}$, illustrating the logistic decay from $\tau_{\max}$ to $\tau_{\min}$.
Figure 8: Illustration of structural fusion: individual causal graphs (left) aggregated into a fused DAG for multi-label event sequences (right) using a simple union.
[…]